OpenStudy (tiffany_rhodes):

Suppose A and B are two events such that P(A) = 0.5 and P(A U B) = 0.8. Find P(B) assuming A and B are independent. I know the answer is 0.6 but I don't understand how to find P(B).

OpenStudy (jtvatsim):

OK, I think I see how we can find it, I'm finishing up the solution.

OpenStudy (tiffany_rhodes):

Alright, thank you :)

OpenStudy (jtvatsim):

We will need two equations. The first is the definition of P(A U B) and the second is \[P(A \cap B)\] when A and B are independent.

OpenStudy (jtvatsim):

When we look these up, we know that \[P(A \cup B) = P(A) + P(B) - P(A \cap B)\] and \[P(A \cap B) = P(A) \times P(B)\] when A and B are independent.

OpenStudy (jtvatsim):

Is that OK so far?

OpenStudy (tiffany_rhodes):

Yeah

OpenStudy (jtvatsim):

Alright, now, what do we know? We know that P(A) = 0.5, and P(A U B) = 0.8. So, when we substitute these values we get this for the first equation: \[0.8 = 0.5 + P(B) - P(A \cap B)\]

OpenStudy (tiffany_rhodes):

Okay, I'm understanding so far.

OpenStudy (jtvatsim):

Great! And this for the second: \[P(A \cap B) = 0.5 \times P(B)\]

OpenStudy (jtvatsim):

Aha! This is the moment of victory! Now, we put the two equations together and get \[0.8 = 0.5 + P(B) - 0.5 P(B)\]

OpenStudy (jtvatsim):

And it follows easily that P(B) = 0.6 by solving... :)

OpenStudy (tiffany_rhodes):

alright, I think I got it. Thanks so much! :)

OpenStudy (jtvatsim):

No problem! Good luck!